OpenStudy (anonymous):

Find the difference. (6x3 + 18x2) - 3x2 =

OpenStudy (anonymous):

\[(6x^3 + 18x^2) - 3x^2=6x^3+(18-3)x^2=6x^3+15x^2\]

OpenStudy (anonymous):

(48+36)-6 =84-6 =78